if y(x) is the solution of dy/dx = (x^(2)+1)/y^(2)

y(0)=2 then y(3)=?
Answer #1

Given D.E. dy dx 1 2 Y(-) = 2 Y(2) is Solution of Di we will find y(3) = ? ye dy = (x²+1) dx Take - Integration Jses = f (x)=) y - x² + 3x + c = ya)= |y) = (x + 3x + c)} Given y (o)= 2 so By Ean D (+360+6) / = 2 2 =) = (2)3 =) 8 Volue Put in . y(x)=